East Texas Energy Inc (ETE) is considering a investing in a new oil reserve. The project will cost $71 million up front and is expected to generate cash flows during each of the following three years, with the amount of cash flow directly related to oil prices. ETE’s finance department has forecast three likely scenarios for oil prices. There is a 10% probability that oil prices will increase, with associated FCF of $40 million per year for three years. There is a 80% probability that oil prices will remain near current levels, with associated FCF of $30 million per year for three years. The third possibility is a 10% probability that oil prices will decline, with associated FCF of $20 million per year for three years. ETE has determined that the required rate of return for this project is 15%. If ETE invests in the new oil reserve, then the investment will also contain an option to expand or grow. This growth option will allow ETE opportunity to replicate the project at the end of the projects original life, with the same cost, FCF, and years of FCF.

Compute the expected NPV of the new oil reserve with the growth option.
Answer #1

In this problem, we need to determine whether to go ahead with investing in the new oil reserve.

Following determinants needs to be calculated to make a decision:

1. Present Value of the Future Cash Inflows in each of the scenarios

2. Multiply each scenarios by the expected probability

3. Calculate NPV of the project and then take a decision

PVAF; (15% , 3 Years ) = 2.28322

PV Factor for year 1, 2 & 3 = 0.8695, 0.7561, 0.6575

Particulars Yearly Cash Flow PVAF PV Cash Flows Probability Multiple
Scenario 1 40 2.28322 91.3288 0.1           9.13
Scenario 2 30 2.28322 68.4966 0.8        54.80
Scenario 3 20 2.28322 45.6644 0.1           4.57
Total        68.50

Other than above, Co has opportunity to replicate the project under growth option.

NPV of the project in a seperate scenario of Growth is 20.3288 (91.3288- 71) with 10% probability. This we expect to receive after 3 years.

PV of replication opportunity = 20.3288* 10% *0.6575 = 1.34

Final NPV of the project= PV Cash Inflow+ PV of replication - Cash Outflow

= 68.50 + 1.34- 71

= - 1.16

As NPV is negative, we should not invest in the project.